# Visible Light Positioning With Lamé Curve LEDs: A Generic Approach for Camera Pose Estimation

### v1

@@ IV-C Experimental Results
In Figs. 88(a) and 8(b), we first keep the camera horizontal and perform positioning at two different heights, i.e., $1.0\,\mathrm{m}$ and $1.4\,\mathrm{m}$, to examine the effect of camera height on the performance of LC-VLP. To more clearly visualize the full 6-DoF camera poses, the top-view camera FoV cone representation in[14] is adopted. It can be observed that, compared with the results at $1.0\,\mathrm{m}$, the estimated camera poses at $1.4\,\mathrm{m}$ are noticeably closer to the reference poses. Specifically, the MPEs at heights of $1.0\,\mathrm{m}$ and $1.4\,\mathrm{m}$ are $5.39\,\mathrm{cm}$ and $3.19\,\mathrm{cm}$, respectively, while the corresponding MREs are $3.64^{\circ}$ and $3.10^{\circ}$. These results indicate that the accuracy of LC-VLP improves as the camera height increases, which is because a higher camera captures larger LED images, which is equivalent to increasing the effective LED size. Enlarging the LED size can lead to higher accuracy, which is consistent with the simulation results. We also note that a lower camera height can increase the likelihood of observing more LEDs, which can partially mitigate the impact of image noise. However, in our experimental setup, the increase in the number of visible LEDs when lowering the camera height is limited; therefore, the LED size becomes the dominant influence factor.
@@ Appendix B Proof of Proposition 1
$\bm{\varpi}(\bm{x}^{\mathrm{c}})=\frac{f}{z^{\mathrm{c}}}\begin{bmatrix}x^{\mathrm{c}}\\ y^{\mathrm{c}}\end{bmatrix}+\begin{bmatrix}u_{0}\\ v_{0}\end{bmatrix}$ (40b)
@@
where
$\left\|\bm{\Xi}\right\|_{2}=\sqrt{\max_{\|\bm{\bm{y}}\|=1}\bm{y}^{\mathsf{T}}\bm{\Xi}\bm{y}}$ (43)
denotes the 2-norm of $\bm{\Xi}$. Let the unit vector $\bm{y}=[y_{1},y_{2}]^{\mathsf{T}}$, then we have:
$\bm{y}^{\mathsf{T}}\bm{\Xi}\bm{y}$ $=y_{1}^{2}+y_{2}^{2}+\left(-y_{1}\cdot\frac{x^{\mathrm{c}}}{z^{\mathrm{c}}}-y_{2}\cdot\frac{y^{\mathrm{c}}}{z^{\mathrm{c}}}\right)^{2}$
$\leq 1+\big(y_{1}^{2}+y_{2}^{2}\big)\left(\left(\frac{x^{\mathrm{c}}}{z^{\mathrm{c}}}\right)^{2}+\left(\frac{y^{\mathrm{c}}}{z^{\mathrm{c}}}\right)^{2}\right)$
$=1+\left(\frac{x^{\mathrm{c}}}{z^{\mathrm{c}}}\right)^{2}+\left(\frac{y^{\mathrm{c}}}{z^{\mathrm{c}}}\right)^{2}.$ (44)
Furthermore, since the LED is captured by the camera, it must lie within the camera’s field of view (FoV). This leads to the following constraints:
@@
which can then be substituted into (43) and (B), resulting in:

### v2

@@ IV-C Experimental Results
In Figs. 88(a) and 8(b), we first keep the camera horizontal and perform positioning at two different heights, i.e., $1.0\,\mathrm{m}$ and $1.4\,\mathrm{m}$, to examine the effect of camera height on the performance of LC-VLP. To more clearly visualize the full 6-DoF camera poses, the top-view camera field of view (FoV) cone representation in[40] is adopted. It can be observed that, compared with the results at $1.0\,\mathrm{m}$, the estimated camera poses at $1.4\,\mathrm{m}$ are noticeably closer to the reference poses. Specifically, the MPEs at heights of $1.0\,\mathrm{m}$ and $1.4\,\mathrm{m}$ are $5.39\,\mathrm{cm}$ and $3.19\,\mathrm{cm}$, respectively, while the corresponding MREs are $3.64^{\circ}$ and $3.10^{\circ}$. These results indicate that the accuracy of LC-VLP improves as the camera height increases, which is because a higher camera captures larger LED images, which is equivalent to increasing the effective LED size. Enlarging the LED size can lead to higher accuracy, which is consistent with the simulation results. We also note that a lower camera height can increase the likelihood of observing more LEDs, which can partially mitigate the impact of image noise. However, in our experimental setup, the increase in the number of visible LEDs when lowering the camera height is limited; therefore, the LED size becomes the dominant influence factor.
@@ Appendix B Proof of Proposition 1
$\bm{\varpi}(\bm{x}^{\mathrm{c}})=\frac{f}{z^{\mathrm{c}}}\begin{bmatrix}x^{\mathrm{c}}\\ y^{\mathrm{c}}\end{bmatrix}+\begin{bmatrix}u_{0}\\ v_{0}\end{bmatrix}$. (40b)
@@
where $\left\|\bm{\Xi}\right\|_{2}=\sqrt{\lambda^{\max}_{(\bm{\Xi}\bm{\Xi}^{\mathsf{T}})}}$ denotes the 2-norm[41] of $\bm{\Xi}$ and $\lambda^{\max}_{(\bm{\Xi}\bm{\Xi}^{\mathsf{T}})}$ is the maximum eigenvalue of $\bm{\Xi}\bm{\Xi}^{\mathsf{T}}$. We first calculate:
$\bm{\Xi}\bm{\Xi}^{\mathsf{T}}=\begin{bmatrix}1+\big(\tfrac{x^{\mathrm{c}}}{z^{\mathrm{c}}}\big)^{2}&\tfrac{x^{\mathrm{c}}}{z^{\mathrm{c}}}\cdot\tfrac{y^{\mathrm{c}}}{z^{\mathrm{c}}}\\ \tfrac{x^{\mathrm{c}}}{z^{\mathrm{c}}}\cdot\tfrac{y^{\mathrm{c}}}{z^{\mathrm{c}}}&1+\big(\tfrac{y^{\mathrm{c}}}{z^{\mathrm{c}}}\big)^{2}\end{bmatrix}$. (43)
Then, let $\det(\bm{\Xi}\bm{\Xi}^{\mathsf{T}}-\lambda\bm{I})=0$ and we obtain the eigenvalues of $\bm{\Xi}\bm{\Xi}^{\mathsf{T}}$ by solving:
$\lambda^{2}-\Big(2+\big(\tfrac{x^{\mathrm{c}}}{z^{\mathrm{c}}}\big)^{2}+\big(\tfrac{y^{\mathrm{c}}}{z^{\mathrm{c}}}\big)^{2}\Big)\lambda+\Big(1+\big(\tfrac{x^{\mathrm{c}}}{z^{\mathrm{c}}}\big)^{2}+\big(\tfrac{y^{\mathrm{c}}}{z^{\mathrm{c}}}\big)^{2}\Big)=0,$ (44)
of which the larger root $\lambda=1+\big(\tfrac{x^{\mathrm{c}}}{z^{\mathrm{c}}}\big)^{2}+\big(\tfrac{y^{\mathrm{c}}}{z^{\mathrm{c}}}\big)^{2}$ is taken as the maximum eigenvalue $\lambda^{\max}_{(\bm{\Xi}\bm{\Xi}^{\mathsf{T}})}$. Therefore, the $\left\|\bm{\Xi}\right\|_{2}$ is calculated as:
$\left\|\bm{\Xi}\right\|_{2}=\sqrt{\lambda^{\max}_{(\bm{\Xi}\bm{\Xi}^{\mathsf{T}})}}=\sqrt{1+\big(\tfrac{x^{\mathrm{c}}}{z^{\mathrm{c}}}\big)^{2}+\big(\tfrac{y^{\mathrm{c}}}{z^{\mathrm{c}}}\big)^{2}}.$ (45)
Furthermore, since the LED is captured by the camera, it must lie within the camera’s FoV. This leads to the following constraints:
@@
which can then be substituted into (45), resulting in: